# USTC-ICTS/PCFT-24-43

### v2

@@ I Introduction
One interesting question regarding holography is to understand how holography could happen in a quantum system. Since the born of AdS/CFT, people have found various models that exhibit holographic features. Besides these explicit constructions, a universal microscopic mechanism that achieves holography is still under exploration. In recent years, people discovered the SYK model [1, 2, 3] as one of the simplest toy models with holographic features. It is a quantum mechanical model of $N$ Majorana fermions in 0+1 dimension with random $q$-local Hamiltonian. Due to its simple structure, we may be able to examine how its holography emerges in much detail.
@@
Another example of a bounded continuous spectrum in the ensemble average is the random matrix theory (e.g. GUE ensemble). If we regard each sample of the random hermitian matrices as the Hamiltonian of some system, the dynamics are clearly chaotic. Indeed, it has been discovered that a double-scaled matrix model is dual to JT gravity [13], which is in spirit similar to the duality between random matrices and string theories [14].
Furthermore, SYK type of models has another appealing feature, namely they have potential experimental realization in various setups [15, 16, 17, 18, 19, 20, 21, 22]. In fact, a very interesting recently proposal [23] claims that on a quantum computer, one can realize a special SYK-like model on a quantum processer and one observe characteristic features of tranversable wormholes in gravity. However, the model in is actually “commuting”, meaning that all terms in the Hamiltonian mutually commute. As a result the model possesses lots of conserved quantities and cannot be quantum chaotic.
The commuting SYK model is further analyzed in [24]. As a random but integrable version of the SYK model, it has an unbounded spectrum of Gaussian distribution. It has been shown in [24] that this integrable variant does not exhibit chaotic behavior though it has some similarities. Interestingly, this provides a new angle to explore the mechanism of quantum chaos (or perhaps holography) in the regular SYK model. Indeed, we can decompose the regular SYK Hamiltonian as the sum over multiple groups, each of which contains only mutually commutating terms and thus is integrable, but different groups do not commute with each other. If we imagine the SYK Hamiltonian as the limit of adding these groups one by one, the chaotic nature emerges as we include huge numbers of these groups, and the integrability is dramatically violated.
@@
$\rho(E)=\mathbb{E}\left[\sum_{\tilde{\lambda}}\delta(E-\tilde{\lambda})\right]$ (2)
@@ II.1 Multi-color chord diagrams
Summing over the last term becomes an involved combinatorial problem, which becomes practically even harder for higher moments because we need to switch multiple pairs of $X^{a}_{I}$. Fortunately, the following double scale limit makes the $n$-th moment computation trackable
@@
For generic $a\neq b$, we show in Appendix A that if $N$ is coprime with any number between 2 and $d-1$, we can always redefine $\psi_{i}$ such that the overlap problem becomes exactly the same as $a=1$ and $b=2$. If the coprime condition does not hold, this equivalence still holds in the large $N$ limit unless $d$ is fine-tuned and comparable with $N$.
@@
This pictorial algorithm generalizes to the computation of any moment $M_{2n}$ (for fixed $n$) because the overlapping of any two neighboring $\tilde{H}_{a}$ with different $a$ in the double scale limit can be regarded as indepedent Poinsson distribution $P(s)$ due to the fact that the total $np$ $\mathcal{X}_{i}^{a}$ in each diagram is sparsely drawn from $N$ $\mathcal{X}^{a}_{i}$’s. Therefore, we summarize the rule as follows:
@@ II.2 Multi-color $q$-harmonic oscillators
We can associate a Fock space representation for the these operators. Denote $\left|0\right\rangle$ as the state for no chords and $\left|n\right\rangle=\alpha^{\dagger n}\left|0\right\rangle$ as the state by creating $n$ open chords. Acting $\alpha$ on $\left|n\right\rangle$ annihilates one of the open chord from the $n$ choices. If we label the open chord from bottom to top in $\left|n\right\rangle$ as 1 to $n$, annihilating the $i$-th open chord will cross $i-1$ open chords below it. Therefore, the total penalty factor for acting $\alpha$ on $\left|n\right\rangle$ is
@@
where $[n]_{q}$ is the $q$ analog of number $n$. After annihilating one open chord, state $\left|n\right\rangle$ becomes proportional to $\left|n-1\right\rangle$
@@
$[\alpha,\alpha^{\dagger}]_{q}\equiv\alpha\alpha^{\dagger}-q\alpha^{\dagger}\alpha=1$ (21)
@@
where the computation is by moving $\alpha$ across all $\alpha^{\dagger}$ using (21) to kill $\left|0\right\rangle$, and $[n]_{q}!=[n]_{q}[n-1]_{q}\cdots[1]_{q}$ is the $q$-factorial. This definition of inner product explicitly treats $\alpha$ as the adjoint of $\alpha^{\dagger}$. It follows that the Hilbert space spanned by $\left|n\right\rangle$ is a well-defined Fock space.
Recall that to compute the $n$-th moment of the SYK Hamiltonian $\tilde{H}_{SYK}$ in the double scale limit, we need to sum over all possible $n$-chord diagrams, which correspond to all Wick contraction patterns of $n$ $\tilde{H}_{SYK}$. Representing a chord diagram on a boundary line as Fig. 2, the contraction of each $\tilde{H}_{SYK}$ can be through either creating an open chord or annihilating an open chord. Therefore, we define a new Hamiltonian $h=\alpha+\alpha^{\dagger}$  [^3] This $h$ is the same as the transfer matrix $T$ in [36, 33] after a similarity transformation. as the avator for the SYK Hamiltonian $\tilde{H}_{SYK}$. In other words, we have
@@
From (24) we see that the spectrum of the double-scaled SYK model is compact and bounded between $-2/\sqrt{1-q}$ and $2/\sqrt{1-q}$. The partition function is given by
[…]
@@
[…]
@@
$[\alpha_{a},\alpha_{a}^{\dagger}]=1,\quad[\alpha_{a},\alpha_{b}^{\dagger}]_{q}=0\quad(a\neq b)$ (33)
@@
=$\mathcal{H}=\bigoplus_{\{n^{a}\}}\mathcal{H}_{\{n^{a}\}}$ (34)
@@ II.3 Double-scaled SYK: $d\rightarrow\infty$ limit
Indeed, the leading order large $d$ correction can be carried out by the same probability crossing argument. For any two crossing chords, they have $1/d$ probability to be the same color and $1-1/d$ probablity to be different colors. Therefore, for each crossing, the averaged panelty factor is
@@ III.2 Spectrum of $H$ for $q=0$
Define a new variable $w=z/d+(1-1/d)/r$, and above equation can be written as
@@ III.2.1 Perturbative in $1/d$
Let us zoom into the edge near $z=-2$ for large but finite $d$ case. It turns out to be convenient to work in $w$ variable rather than $r$. Writing the LHS of (70) as $(1-1/d)/(w-z/d)$, we can decompose (70) into two equations respectively for the real and imaginary parts
@@ III.2.2 Non-perturbative in $1/d$
As these expressions have singularity at $z=-2$, they are valid only for $z<-2$. Taking them into (85), expanding it in large negative $y=z+2$ leads to
@@ IV.1 Coarse-grained path integral
$Z(\beta)$ =$=\sum_{\{n_{ij}^{a}\}}q^{\sum_{a\neq b}\sum_{i>k>j>l}n_{ij}^{a}n_{kl}^{b}}$
×$\times\sum_{\{{\bf a}_{i}\}}\prod_{i=1}^{s}\left(\mathcal{C}(n_{ij}^{a};{\bf a}_{i};q)\frac{\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle}{\left\langle{\bf a}_{i}|{\bf a}_{i}\right\rangle\dim\mathcal{H}_{\{n^{a}_{i}\}}}\right)$ (93)
@@
In (93), we split $\beta$ Euclidean time into $s$ segments with length $\beta/s$. The expectation value $\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle$ represents the first step of generating $\{n^{a}_{i}\}$ open chords. The function $\mathcal{C}(n_{ij}^{a}$;${\bf a}_{i};q)$ depending on $n_{ij}^{a}$ with $j\neq i$ and includes the contribution from splitting and reordering of $n_{i}$ open chords.
To solve $\mathcal{C}$ and $\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle$ in closed form is very hard. Nevertheless, for $q$ close to 1 and Schwarzian limit, as argued in [26], we can replace them with the known result for $q=1$ but keep $q$ nontrivial in the first line of (93). At $q=1$ the algebra of (33) becomes $d$ independent harmonic oscillators, and all states $\left|{\bf a}_{i}\right\rangle$ with the same $\{n^{a}_{i}\}$ is equivalent to the number state |$\left|\{n^{a}_{i}\}\right\rangle=\prod_{a}(\alpha_{a}^{\dagger})^{n^{a}_{i}}\left|0\right\rangle$.
Summing over $\{{\bf a}_{i}\}$ for a fixed $\{n^{a}_{i}\}$ gives the dimension of the subspace $\dim\mathcal{H}_{\{n^{a}_{i}\}}$, which cancels the same term in the denominator of (93). Given $\{n_{ij}^{a}\}$ and $\{n^{a}_{i}\}$, there are $\prod_{a}(n^{a}_{i}!/\prod_{j\neq i}(n_{ij}^{a})!)$ ways to split the $n_{i}$ open chords into $(s-1)d$ groups. Moreover, for fixed $a,i,j$, the $n_{ij}^{a}$ color $a$ open chords can self intersect and give a contribution $\prod_{i>j}\prod_{a}(n_{ij}^{a})!$. This is the reordering process. Lastly, for $d$ harmonic oscillators, we can easily compute
@@
[…]
@@
[…]
@@
[…]
@@ IV.3 The quadratic action
Since $K$ is a symmetric real matrix, it can be diagonalized by an orthonormal matrix. The eigenvalues of $K$ include one $d-1$ and $d-1$ numbers of $-1$ with multiplicity $d-1$. Along the diagonal basis of $\delta J^{a}$, we can split it into $\delta J$ and $\delta J^{i}$ and rewrite the quadratic action as
@@
±=$\mathcal{M}_{\pm}=\{n_{\eta}|P_{(n_{\eta}/\omega-1)/2}^{1/2-h}(\cos\theta_{r})\pm P_{(n_{\eta}/\omega-1)/2}^{1/2-h}(-\cos\theta_{r})=0\}$ (125)
Note that $n_{\eta}\rightarrow-n_{\eta}$ gives the same solution because of the symmetry $P_{\nu}^{\mu}(z)=P_{-1-\nu}^{\mu}(z)$. The independent solutions are labeled by the eigenvalues ±$\mathcal{M}_{\pm}/\mathbb{Z}_{2}$, where $\mathbb{Z}_{2}$ refers to this reflection symmetry. On the other hand, since $\tau$ labels the segments on a circle, we need to impose some periodic boundary condition for $\tau$. In $(\tau_{1},\tau_{2})$ coordinate, we should require $f(\beta_{r},\tau_{2})=f(\tau_{2},0)$, which in $(\tau,\theta)$ coordinate means
@@
[…]
@@ IV.4.1 $h=2$
$\mathcal{M}_{+}$ =$=\{n\neq 0|\tan\frac{x}{2}\tan\frac{n}{4}+\frac{n}{2x}=0\}$ (137)
$\mathcal{M}_{-}$ =$=\{|n|\neq 2x|\tan\frac{x}{2}\cot\frac{n}{4}-\frac{n}{2x}=0\}$ (138)
@@
[…]

### v4

@@ I Introduction
One interesting question regarding holography is understanding how holographic duality could emerge in quantum systems. Since the birth of AdS/CFT correspondence, researchers have identified various models exhibiting holographic features. Beyond these explicit constructions, a universal microscopic mechanism underlying holography remains under exploration. In recent years, the SYK model [1, 2, 3] has emerged as one of the simplest toy models possessing holographic properties. This quantum mechanical model describes $N$ Majorana fermions in (0+1) dimensions with a random $q$-local Hamiltonian. Owing to its simplified structure, it allows detailed examination of holographic emergence.
@@
Another example of a bounded continuous spectrum in the ensemble average appears in random matrix theory (e.g., the GUE ensemble). Treating each random Hermitian matrix sample as as the Hamiltonian of some system, the system exhibits explicitly chaotic dynamics. Notably, a double-scaled matrix model has been shown dual to JT gravity [13], which is in spirit analogous to the duality between random matrices and string theories [14].
Furthermore, SYK type models possess another appealing feature: potential experimental realizations in various platforms [15, 16, 17, 18, 19, 20, 21, 22]. Intriguingly, a recent proposal [23] suggests that certain SYK-like models implemented on quantum processors could exhibit signatures of traversable wormholes. However, the specific model proposed is ”commuting” – all Hamiltonian terms mutually commute – resulting in numerous conserved quantities that preclude quantum chaos.
The commuting SYK model is further analyzed in [24]. As a random yet integrable SYK variant, it displays an unbounded Gaussian-distributed spectrum. It has been demonstrated in [24] that this integrable system does not exhibit chaotic behavior despite superficial similarities. Interestingly, this provides a new angle to explore the mechanism of quantum chaos (or perhaps holography) in the regular SYK model. Specifically, we can decompose the regular SYK Hamiltonian into multiple groups of mutually commuting terms (and is thus integrable), but different groups do not commute with each other. If we imagine the SYK Hamiltonian as the limit of adding these groups one by one, the chaotic nature emerges as we include huge numbers of these groups, and the integrability is dramatically violated.
@@
[…]
@@ II.1 Multi-color chord diagrams
Summing over the last term becomes an involved combinatorial problem, which becomes practically even harder for higher moments because we need to switch multiple pairs of $X^{a}_{I}$. Fortunately, the following double-scaling limit renders the computation of the $n$-th moment tractable
@@
For generic $a\neq b$, we show in Appendix A that if $N$ is coprime with any number between 2 and d1d-1, we can always redefine $\psi_{i}$ such that the overlap problem becomes exactly the same as $a=1$ and $b=2$. If the coprime condition does not hold, this equivalence still holds in the large $N$ limit unless $d$ is fine-tuned and comparable with $N$.
@@
This pictorial algorithm generalizes to the computation of any moment $M_{2n}$ (for fixed $n$) because the overlapping of any two neighboring $\tilde{H}_{a}$ with different $a$ in the double scale limit can be regarded as indepedent Poisson distribution $P(s)$ due to the fact that the total npnp $\mathcal{X}_{i}^{a}$ in each diagram is sparsely drawn from NN $\mathcal{X}^{a}_{i}$’s. Therefore, we summarize the rule as follows:
@@ II.2 Multi-color $q$-harmonic oscillators
We can associate a Fock space representation for the these operators. Denote $\left|0\right\rangle$ as the state for no chords and $\left|n\right\rangle=\alpha^{\dagger n}\left|0\right\rangle$ as the state by creating $n$ open chords. Acting $\alpha$ on $\left|n\right\rangle$ annihilates one of the open chord from the $n$ choices. If we label the open chord from bottom to top in $\left|n\right\rangle$ as 1 to $n$, annihilating the $i$-th open chord will cross i1i-1 open chords below it. Therefore, the total penalty factor for acting $\alpha$ on $\left|n\right\rangle$ is
@@
where $n]_{q}$ is the $q$ analog of number $n$. After annihilating one open chord, state $\left|n\right\rangle$ becomes proportional to $\left|n-1\right\rangle$
@@
$\alpha,\alpha^{\dagger}]_{q}\equiv\alpha\alpha^{\dagger}-q\alpha^{\dagger}\alpha=1$ (21)
@@
where the computation is by moving $\alpha$ across all $\alpha^{\dagger}$ using (21) to kill $\left|0\right\rangle$, and $n]_{q}!=[n]_{q}[n-1]_{q}\cdots[1]_{q}$ is the $q$-factorial. This definition of inner product explicitly treats $\alpha$ as the adjoint of $\alpha^{\dagger}$. It follows that the Hilbert space spanned by $\left|n\right\rangle$ is a well-defined Fock space.
Recall that to compute the $n$-th moment of the SYK Hamiltonian $\tilde{H}_{SYK}$ in the double scale limit, we need to sum over all possible $n$-chord diagrams, which correspond to all Wick contraction patterns of $n$ $\tilde{H}_{SYK}$. Representing a chord diagram on a boundary line as Fig. 2, the contraction of each $\tilde{H}_{SYK}$ can be through either creating an open chord or annihilating an open chord. Therefore, we define a new Hamiltonian $h=\alpha+\alpha^{\dagger}$  [^3] This $h$ is the same as the transfer matrix $T$ in [36, 33] after a similarity transformation. as the avatar for the SYK Hamiltonian $\tilde{H}_{SYK}$. In other words, we have
@@
From (24) we see that the spectrum of the double-scaled SYK model is compact and bounded between $2/\sqrt{1-q}$ and $2/\sqrt{1-q}$. The partition function is given by
$Z(\beta)=2^{N}\left\langle 0|e^{-\beta h}|0\right\rangle=2^{N}\int_{0}^{\pi}d\theta\mu(\theta)e^{-\beta E(\theta)}$ (29)
@@
where $E_{0}=-2/\sqrt{1-q}$ is the ground energy, and we have rescaled the normalization such that $\int dE\rho(E)=1$.
@@
$\alpha_{a},\alpha_{a}^{\dagger}]=1,\quad[\alpha_{a},\alpha_{b}^{\dagger}]_{q}=0\quad(a\neq b)$ (33)
@@
$\mathcal{H}=\bigoplus_{\{n^{a}\}}\mathcal{H}_{\{n^{a}\}}$ (34)
@@ II.3 Double-scaled SYK: $d\rightarrow\infty$ limit
Indeed, the leading order large $d$ correction can be carried out by the same probability crossing argument. For any two crossing chords, they have $1/d$ probability to be the same color and 11/d1-1/d probablity to be different colors. Therefore, for each crossing, the averaged penalty factor is
@@ III.2 Spectrum of $H$ for $q=0$
Define a new variable w=z/d+(11/d)/rw=z/d+(1-1/d)/r, and above equation can be written as
@@ III.2.1 Perturbative in $1/d$
Let us zoom into the edge near z=2z=-2 for large but finite $d$ case. It turns out to be convenient to work in $w$ variable rather than $r$. Writing the LHS of (70) as (11/d)/(wz/d)(1-1/d)/(w-z/d), we can decompose (70) into two equations respectively for the real and imaginary parts
@@ III.2.2 Non-perturbative in $1/d$
As these expressions have singularity at z=2z=-2, they are valid only for z<2z<-2. Taking them into (85), expanding it in large negative $y=z+2$ leads to
@@ IV.1 Coarse-grained path integral
$Z(\beta)$ $=\sum_{\{n_{ij}^{a}\}}q^{\sum_{a\neq b}\sum_{i>k>j>l}n_{ij}^{a}n_{kl}^{b}}$
$\times\sum_{\{{\bf a}_{i}\}}\prod_{i=1}^{s}\left(\mathcal{C}(n_{ij}^{a};{\bf a}_{i};q)\frac{\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle}{\left\langle{\bf a}_{i}|{\bf a}_{i}\right\rangle\dim\mathcal{H}_{\{n^{a}_{i}\}}}\right)$ (93)
@@
In (93), we split $\beta$ Euclidean time into $s$ segments with length $\beta/s$. The expectation value $\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle$ represents the first step of generating $\{n^{a}_{i}\}$ open chords. The function $\mathcal{C}(n_{ij}^{a}$${\bf a}_{i};q)$ depending on $n_{ij}^{a}$ with $j\neq i$ and includes the contribution from splitting and reordering of $n_{i}$ open chords.
To solve $\mathcal{C}$ and $\left\langle{\bf a}_{i}|e^{-\beta H/s}|0\right\rangle$ in closed form is very hard. Nevertheless, for $q$ close to 1 and Schwarzian limit, as argued in [26], we can replace them with the known result for $q=1$ but keep $q$ nontrivial in the first line of (93). At $q=1$ the algebra of (33) becomes $d$ independent harmonic oscillators, and all states $\left|{\bf a}_{i}\right\rangle$ with the same $\{n^{a}_{i}\}$ is equivalent to the number state $\left|\{n^{a}_{i}\}\right\rangle=\prod_{a}(\alpha_{a}^{\dagger})^{n^{a}_{i}}\left|0\right\rangle$.
Summing over $\{{\bf a}_{i}\}$ for a fixed $\{n^{a}_{i}\}$ gives the dimension of the subspace $\dim\mathcal{H}_{\{n^{a}_{i}\}}$, which cancels the same term in the denominator of (93). Given $\{n_{ij}^{a}\}$ and $\{n^{a}_{i}\}$, there are $\prod_{a}(n^{a}_{i}!/\prod_{j\neq i}(n_{ij}^{a})!)$ ways to split the $n_{i}$ open chords into (s1)d(s-1)d groups. Moreover, for fixed $a,i,j$, the $n_{ij}^{a}$ color $a$ open chords can self intersect and give a contribution $\prod_{i>j}\prod_{a}(n_{ij}^{a})!$. This is the reordering process. Lastly, for $d$ harmonic oscillators, we can easily compute
@@
$=$ $\frac{1}{\sqrt{\det(-2\pi\lambda M)}}\int d{\bf J}e^{{\bf n}\cdot{\bf J}+\frac{1}{2\lambda}{\bf J}\cdot M^{-1}\cdot{\bf J}}$ (99)
@@
$Z(\beta)=\int d{\bf J}\frac{e^{\frac{1}{2\lambda}{\bf J}\cdot M^{-1}\cdot{\bf J}}}{\sqrt{\det(-2\pi\lambda M)}}e^{\left(\frac{1}{s}+\frac{2}{s^{2}d}\sum_{i>j;a}e^{J_{ij}^{a}}\right)\beta^{2}/2}$ (100)
@@
and the partition function is $Z(\beta)\sim\int d{\bf J}e^{S/\lambda}$. In $q\rightarrow 1$ limit, we have $\lambda\sim 1-q\ll 1$. Therefore, the partition function can be computed by saddle approximation. Variation of $\bf J$} leads to
@@ IV.3 The quadratic action
Since $K$ is a symmetric real matrix, it can be diagonalized by an orthonormal matrix. The eigenvalues of $K$ include one d1d-1 and d1d-1 numbers of 1-1 with multiplicity d1d-1. Along the diagonal basis of $\delta J^{a}$, we can split it into $\delta J$ and $\delta J^{i}$ and rewrite the quadratic action as
@@
$\mathcal{M}_{\pm}=\{n_{\eta}|P_{(n_{\eta}/\omega-1)/2}^{1/2-h}(\cos\theta_{r})\pm P_{(n_{\eta}/\omega-1)/2}^{1/2-h}(-\cos\theta_{r})=0\}$ (125)
Note that $n_{\eta}\rightarrow-n_{\eta}$ gives the same solution because of the symmetry $P_{\nu}^{\mu}(z)=P_{-1-\nu}^{\mu}(z)$. The independent solutions are labeled by the eigenvalues $\mathcal{M}_{\pm}/\mathbb{Z}_{2}$, where $\mathbb{Z}_{2}$ refers to this reflection symmetry. On the other hand, since $\tau$ labels the segments on a circle, we need to impose some periodic boundary condition for $\tau$. In $(\tau_{1},\tau_{2})$ coordinate, we should require $f(\beta_{r},\tau_{2})=f(\tau_{2},0)$, which in $(\tau,\theta)$ coordinate means
@@
$=$ $\exp\left[\frac{1}{4\pi i}\int_{\mathcal{C}_{\pm}}dz\log G_{\pm}(z)\frac{F_{\pm}^{\prime}(z)}{F_{\pm}(z)}\right]$ (134)
@@ IV.4.1 $h=2$
$\mathcal{M}_{+}$ $=\{n\neq 0|\tan\frac{x}{2}\tan\frac{n}{4}+\frac{n}{2x}=0\}$ (137)
$\mathcal{M}_{-}$ $=\{|n|\neq 2x|\tan\frac{x}{2}\cot\frac{n}{4}-\frac{n}{2x}=0\}$ (138)
@@
$I_{\pm}=-\frac{1}{4\pi i}\int_{0}^{4x^{2}\tan^{2}x/2}d\alpha\int_{\mathcal{C}_{\pm}}dz\frac{1}{\alpha+z^{2}}\frac{F_{\pm}^{\prime}(z)}{F_{\pm}(z)}$ (143)